40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun May 16 12:51:08 2010, kiwamu, Update, Green Locking, reflected beam at PD 
    Reply  Sun May 16 19:25:45 2010, Koji, Update, Green Locking, reflected beam at PD 
       Reply  Thu May 20 13:29:40 2010, kiwamu, Update, Green Locking, mode profile at PPKTP crystal PPKTPmode.png
          Reply  Thu May 20 14:18:59 2010, kiwamu, Update, Green Locking, mode profile of 40m cavity MMT40mcavity.png
             Reply  Wed May 26 04:14:21 2010, kiwamu, Update, Green Locking, locked OLTF_endPDH.png
                Reply  Wed May 26 14:38:02 2010, Koji, Update, Green Locking, locked 
                   Reply  Thu May 27 02:22:24 2010, kiwamu, Update, Green Locking, more details stimes.pngspectrum.pngETMX_small.png
                      Reply  Thu May 27 09:43:50 2010, rana, Update, Green Locking, more details 
                         Reply  Mon Jul 12 13:16:53 2010, kiwamu, Update, Green Locking, PZT feedback at X end arm_day_night.png
                      Reply  Wed Jun 9 11:13:11 2010, kiwamu, Update, Green Locking, lock with PDH box 
                         Reply  Fri Jun 11 11:54:42 2010, kiwamu, Update, Green Locking, end PDH with thermal feedback LPF.pngthermal_feedback.png
                            Reply  Fri Jun 11 13:32:28 2010, Koji, Update, Green Locking, end PDH with thermal feedback 
                         Reply  Sun Jun 20 18:28:25 2010, kiwamu, Update, Green Locking, Re: lock with PDH box 
                Reply  Wed Jul 21 19:11:18 2010, Dmass, Update, Green Locking, locked 
Message ID: 3065     Entry time: Fri Jun 11 11:54:42 2010     In reply to: 3059     Reply to this: 3066
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: end PDH with thermal feedback 

A thermal feedback was installed to the end PDH locking and it works well. There are no saturations 

As I said the feedback signal was sometimes saturated at the sum-amp because the drive signal going to the laser PZT was large at low frequency (below 1Hz).

So I made a passive low pass filter which filters the signal controlling the temperature of the laser crystal, and put it before the temperature drive input.

Now the amount of the feedback signal got reduced when it is locked, and there are no saturations. It's very good.


(thermal property of the crystal) 

According to the specification sheet for the 1W Innolight, the thermal properties of the crystal are:

  Response coefficient : 3GHz/K 

  Temperature control coefficient : 1K/V

  Thermal response bandwidth: 1Hz 

 

(filter circuit and actuator response)

In order to feedback the signal blow 1Hz, a low pass fiter is needed. 

The attachment:1 shows the filter circuit I made.

Since I found that the drive input had an input impedance of 100kOhm, so I put relatively big resistors to have a moderate gain.

The expected actuator responses are also attached.

The blue curve represents the response of the PZT, the green is the thermal response including the low pass filter and the red curve is the total response composed of both the responses.

I assume that the PZT response is 1MHz/V according to Mott's measurement.

Also I assume that the thermal response intrinsically has two poles at 1Hz according to the specification listed above.

In the total response, there is a little gain reduction around 2Hz due to the cancelation of each other, but it still looks okay.

 

Attachment 1: LPF.png  37 kB  | Show | Hide all | Show all
Attachment 2: thermal_feedback.png  60 kB  | Hide | Hide all | Show all
thermal_feedback.png
ELOG V3.1.3-